OpenStudy (mattduhhelper):

The product of 5 and a number x is 1/4 . What is the value of x? x=−194x=−194 x=120x=120 x=54x=54 x=434

OpenStudy (mattduhhelper):

@Awolflover1

OpenStudy (tteesstt8080):

Here: Pretend we in Mathlandia. no words. *cheer or groan* you see x, 5, and 1/4. don't question x, it's VIP lemme sneak in some words product of 5 AND a number x IS 1/4 translate into numbers? 5x=1/4 Solve for x x=1/4 *1/5 x=1/20

OpenStudy (tteesstt8080):

now those answer choices look fishy yet this is what I did

OpenStudy (mathmale):

Let the number be x. Then "the product of 5 and a number" comes out to 5x. 5x = 1/4 Can you solve this for x?